A pipe closed at one end is sounding its fundamental note. If the length of the pipe is 10cm, calculate the wavelength of the note. (neglect end correction of the pipe).
1
Expert's answer
2020-09-10T14:40:18-0400

The fundamental frequency (with "n = 1" ) in a pipe closed at one end is given by the following expression:


"f = \dfrac{nv}{4L} = \dfrac{v}{4L}"

where "v" is the speed of sound in air, "L = 10cm = 0.1m" is the length of the pipe. The wavelength will be:


"\lambda = \dfrac{v}{f} = 4L = 4\cdot 0.1m = 0.4m"

Answer. 0.4 m.